OpenStudy (anonymous):

Find the power series representation centered at x=0 of the following function: F(x) = 1/(1+2(x^2))

OpenStudy (anonymous):

easy way may be to start with \[\frac{1}{1-x}=1+x+x^2+...=\sum_{k=0}^{\infty}x^k\] and rework \[\frac{1}{1+2x^2}\] so look like that

OpenStudy (anonymous):

\[\frac{1}{1+x}=1-x+x^2-x^3+...=\sum_{k=0}^{\infty}(-1)^kx^k\] \[\frac{1}{1+x^2}=1-x^2+x^4-...=\sum_{k=0}^{\infty}(-1)^kx^{2k}\] one more and you are done

OpenStudy (anonymous):

so it should look something like this then?? \[\sum_{n=0}^{\infty } (-1)^{n} (1/2n)(x^{2n}) \]

OpenStudy (anonymous):

or maybe that (1/2n) should be 1/(n+1)